OpenStudy (anonymous):

solve the following for x: 3e^(-0.15x+1)=6

OpenStudy (anonymous):

First isolate the \(e^{-.15x + 1}\)

OpenStudy (anonymous):

So divide 3 out?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

leaves you with e^(-0.15x+1)=(6/3) and I'm still stuck...

OpenStudy (anonymous):

Well 6/3 = 2. Then you take the natural log of both sides

OpenStudy (anonymous):

Remember that \[ln\ e^a = a\]

OpenStudy (anonymous):

obviously correct me if I'm wrong but would that leave me with -0.15x+1/x-2/x=0?

OpenStudy (anonymous):

I don't think so. \[3e^{(-0.15x+1)}=6\] \[\implies e^{-.15x + 1} = 2\] \[\implies ln\ e^{-.15x + 1} = ln 2\] \[\implies -.15x + 1 = ln 2\] \[\implies -.15x = (ln2)-1\] \[\implies x = \frac{(ln2) -1}{-.15}\]

OpenStudy (anonymous):

Oh. Hmm what if you have a polynomial. The example I have is 2^(2x)+2^(x)-6=0. Are the steps much different?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

For that you'd want to start by solving the quadratic.

OpenStudy (anonymous):

Let \(k = 2^x \implies 2^{2x} = (2^x)^2 = k^2\) \[\implies 2^{2x} + 2^x -6 = k^2 + k -6 = 0\] Use the quadratic formula (or factoring) to solve for k.

OpenStudy (anonymous):

0=-3, 2?

OpenStudy (anonymous):

k = -3 and k = 2.

OpenStudy (anonymous):

whoops that's what I meant!

OpenStudy (anonymous):

Since k = \(2^x\) that means that \(2^x = -3\) and \(2^x = 2\) are solutions.

OpenStudy (anonymous):

So to find x we again take the log of both sides.

OpenStudy (anonymous):

from which equation?

OpenStudy (anonymous):

From those last two equations.

OpenStudy (anonymous):

\[2^x = -3\] and \[2^x = 2\]

OpenStudy (anonymous):

Sorry for the extremely late reply wouldn't that lead you with ln2^x=-3ln and ln2^x=2ln? but to get x alone I'm lost

OpenStudy (anonymous):

Ah, well when you take the log of something you can bring out (and down) the exponent. \[ln\ 2^x= x(ln\ 2)\]

OpenStudy (anonymous):

So you have \[x(ln\ 2) = ln\ 2\] and \[x(ln\ 2) = ln\ -3\] But (ln -3) is not defined, so the first solution is the only valid one. And when you divide both sides by (ln 2) you have x = 1.